OpenStudy (anonymous):

Multiply or divide as indicated. (a+b)^9/(a+b)^4

OpenStudy (anonymous):

In general \[\frac{n^{x}}{n^{y}} = n^{x-y}\]

OpenStudy (anonymous):

sorry i dont really get it

OpenStudy (anonymous):

If you're dividing the same mantissa (i.e., the thing you're raising to an exponent) with different exponents, you can just subtract the exponents.

OpenStudy (vjordillo):

If they have the same base,n , and their exponents are different, just copy the base and subtract the exponent of divisor from the exponent of dividend.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

what about this one? 10x^5/2x^2

OpenStudy (vjordillo):

They have the same base, x, but differs in their coefficients. First thing to do is to disregard for the meantime their coefficients. Then work just like the other. After getting the answer for the variable, multiple it to the quotient of their coefficients.

OpenStudy (anonymous):

can you show an example?

OpenStudy (vjordillo):

\[\[\frac{ 6x ^{5} }{ 4x ^{2} }=(\frac{ 6 }{ 4 })(x ^{5-2})=(\frac{ 6 }{ 4 })(x ^{3})=(\frac{ 3 }{ 2 })(x ^{3})=\frac{ 3 }{ 2 }x ^{3}\]\]